### v1

# SequencePAR: Understanding Pedestrian Attributes via A Sequence Generation Paradigm

Abstract
Current pedestrian attribute recognition (PAR) algorithms are developed based on multi-label or multi-task learning frameworks, which aim to discriminate the attributes using specific classification heads. However, these discriminative models are easily influenced by imbalanced data or noisy samples. Inspired by the success of generative models, we rethink the pedestrian attribute recognition scheme and believe the generative models may perform better on modeling dependencies and complexity between human attributes. In this paper, we propose a novel sequence generation paradigm for pedestrian attribute recognition, termed SequencePAR. It extracts the pedestrian features using a pre-trained CLIP model and embeds the attribute set into query tokens under the guidance of text prompts. Then, a Transformer decoder is proposed to generate the human attributes by incorporating the visual features and attribute query tokens. The masked multi-head attention layer is introduced into the decoder module to prevent the model from remembering the next attribute while making attribute predictions during training. Extensive experiments on multiple widely used pedestrian attribute recognition datasets fully validated the effectiveness of our proposed SequencePAR. The source code and pre-trained models will be released at https://github.com/Event-AHU/OpenPAR.
Index Terms:
Pedestrian Attribute Recognition, Pre-trained Big Models, Prompt Learning, Image Captioning, Sequence Generation
I Introduction
Pedestrian Attribute Recognition (PAR) [1, 2] involves utilizing a predefined set of attributes to estimate or infer semantic characteristics for pedestrian images, such as age, gender, dress, movement, etc. It plays an important role in many practical scenarios like intelligent video monitoring, autonomous driving, and pedestrian analysis. Pedestrian attribute recognition also plays a significant role in other pedestrian-related tasks, such as assisting in matching IDs for pedestrian re-identification tasks [3, 4, 5], object detection [6], natural language description-based pedestrian search or retrieval tasks [7, 8, 9], and analyzing and mining pedestrian attributes to predict pedestrian behavior intentions and possible action trajectories [10].
Figure 1: Illustration of different pedestrian attribute recognition frameworks based on Transformer networks. T.E. and L.E. are short for Transformer Encoder and Language Encoder, respectively. Note that existing methods (a, b) follow the discriminative framework, while ours (c) belongs to the generative paradigm.
Due to its irreplaceable important role, many deep learning based pedestrian attribute recognition algorithms have been proposed recently and made great progress. Specifically, numerous researchers have utilized Convolutional Neural Networks (CNNs) to acquire feature representations from pedestrian images. Subsequently, they employ fully connected layers [11, 12] to regress the responses of each attribute. To model the relations between human attributes, the Recurrent Neural Network (RNN) and the Graph Neural Network (GNN) are introduced to further enhance the CNN features [13, 14, 15]. Recently, the Transformer [16, 17] is also adapted for pedestrian attribute recognition [18, 19, 20, 21, 22] due to its strong ability on modeling the long-range global relations. For example, Cheng et al. propose the VTB [18] which achieves high-performance attribute recognition via multimodal Transformer based image-text fusion. The introduction of these algorithms has greatly promoted the development of the pedestrian attribute recognition field.
However, we find that these pedestrian attribute recognition algorithms basically adopt the framework of multi-label classification or multi-task learning, which belongs to the category of discriminative models. It is well known that these discriminative models are susceptible to the following factors:
  • Existing works directly utilize the CNNs/Transformers to predict attribute response scores which can be easily disturbed by the issue of data imbalance. A large number of negative samples may also lead to sparse attribute prediction due to the issue of data sparsity.
  • The annotation of human attributes is a challenging task (for example, some datasets [23] provide uncertain attribute annotations), and the annotated labels for the training data may contain noise or errors, which can affect the performance and reliability of the model.
  • The semantic connection between attributes is weak due to the simultaneous regression of each attribute.
As we all know, generative models can better model dependencies and complexity between data over discriminative models. It describes the relationships between multiple tasks through joint probability distributions and thus performs well in capturing the complexity and diversity of human attributes. The generative models can improve model performance and generalization by incorporating external knowledge or prior information. Therefore, a highly intuitive question worthy of consideration is: “Can we break away from the category of discriminative models and re-examine pedestrian attribute recognition from the perspective of generative models?"
Inspired by the aforementioned observations and thinking, in this work, we propose a novel sequence generation paradigm for pedestrian attribute recognition, termed SequencePAR. A comparison between existing Transformer based pedestrian attribute recognition algorithms and our SequencePAR is illustrated in Fig. 1. The key insight of this work is that we formulate attribute recognition as an image captioning task, which can model the relations between human attributes by generating descriptive captions for human images. This generative model allows us to encapsulate the underlying connections between attributes. Furthermore, the burgeoning evidence from various large-scale models strongly indicates the exceptional performance of big language models in precisely this domain. In our practical implementation, given the pedestrian image, we first partition it into non-overlapping tokens and adopt the CLIP visual encoder [24] to get the visual tokens due to its strong feature expression ability and generalization. We also take the attribute descriptions as the input to capture the high-level semantic information. We adopt the word embedding to transform the attributes into text embeddings and concatenate them with text prompts as the attribute query token. Then, we feed the attribute query tokens into the masked multi-head attention layer, and the outputs are fed into the multi-head attention layer together with visual tokens for attribute generation. Note that, the normalization layers and feed-forward layers are also incorporated in the sequence generation decoder network. In the testing phase, the greedy search mechanism is adopted for attribute generation. An overview of our proposed SequencePAR framework can be found in Fig. 2.
To sum up, the key contributions of this paper can be summarized as the following three aspects:
  • $\bullet$
    We propose a new generative pedestrian attribute recognition framework, termed SequencePAR, which broke away from the fixed mindset of multi-label classification. It is the first work to handle PAR in a sequence generation manner which handles the issue of imbalanced and noisy attribute learning to some extent.
  • $\bullet$
    We propose a novel masked Transformer decoder that predicts each attribute sequentially based on pedestrian tokens and textual representations. It addresses the issue of weak connection of attribute context in the standard multi-label classification.
  • $\bullet$
    Extensive experiments on multiple popular pedestrian attribute recognition datasets fully validated the effectiveness of our proposed SequencePAR.
The rest of this paper is organized as follows: We review related works most related to our paper in Section II. The approach we proposed is mainly introduced in Section III, including the problem formulation, overview, input representation, sequence generation module, and loss function. In Section IV, we conduct extensive experiments and give an in-depth analysis of multiple benchmark datasets. We conclude this paper and propose possible research directions in Section V.
II Related Work
In this section, we will give a brief introduction to the related works on Pedestrian Attribute Recognition, Transformer Networks and Sequence Generation. More related works can be found in the following survey [1] 111github.com/wangxiao5791509/Pedestrian-Attribute-Recognition-Paper-List.
II-A Pedestrian Attribute Recognition
Pedestrian attribute recognition methods can be categorized into various approaches, such as CNN, RNN, graphic model-based algorithms, attention mechanisms, and Transformer-based methods. Early CNN-based approaches have demonstrated promising results in pedestrian attribute recognition. For instance, MTCNN [11] proposed a multi-task learning approach, utilizing multiple CNNs to learn attribute-specific features while sharing knowledge among them. PGDM [25] predicts key points of the human body to leverage deformable body structure information for pedestrian attribute prediction, which captures pose variations and enhances attribute prediction performance. The RNN model is introduced into the PAR community to model the semantic association between human attributes. To be specific, Wang et al. [26] employed LSTM to establish high-level semantic dependencies between labels in pedestrian attribute recognition. By incorporating the previously predicted labels, visual features can be dynamically adjusted for the subsequent moments. Zhao et al. [27] employ convolutional LSTM to explore correlations between attribute groups. They also use recursive attention to leverage spatial locality within each group and attentional correlations between different attribute groups. Graphic models can be employed to establish connections or relationships between attributes, as well as between attributes and pedestrian components. Specifically, Chen et al. [28] integrate multiple features by combining them and feeding them into a classifier to obtain confidence levels. Subsequently, they employed conditional random fields (CRF) to explore and extract associations between attributes. The attention mechanism is widely used in deep neural networks and attribute recognition tasks. For example, Liu et al. introduce HydraPlus-Net [29], an architecture that leverages the inception-v2 [30] framework to generate multi-level feature maps. This approach enables the network to capture diverse pedestrian details and extract more comprehensive features.
Inspired by the success of Transformer networks in natural language processing and computer vision tasks, some researchers also adapted Transformers [16, 17, 31] for pedestrian attribute recognition. For instance, Weng et al. [32] utilize the Attribute Location Module(ALM) to obtain attribute-specific feature representations, and leverage the Transformer to enhance the connection between attribute-related features and the modeled attributes. Cheng et al. [18] approached pedestrian attribute recognition as a visual language task by converting the attribute list into text features. They fed both the visual features extracted from pedestrian images and the text features into a Transformer model, facilitating effective interaction between the modalities. Different from these works, we employ a generative framework that formulates pedestrian attribute recognition as a sequence generation problem. Through iterative attribute generation, our approach explicitly models the semantic relationships between attributes. This addresses the limitations of traditional pedestrian attribute recognition methods, which often overlook the modeling of relationships within attributes or rely on manual assumptions.
II-B Transformer Network
Based on self-attention mechanisms, the Transformer [16] network first caused a significant stir in the field of natural language processing, demonstrating performance far superior to RNN/LSTM [33] series algorithms across various text-related tasks. Then, the pre-training techniques developed based on the Transformer further demonstrate its effectiveness, for example, the BERT [31], GPT series [34, 35], LLaMA [36], LaMDA [37], Baichuan-2 [38]. The remarkable success of the Transformer in natural language processing (NLP) has inspired researchers to explore its application in computer vision. Significant efforts have been made to introduce the Transformer into multiple fields. For example, the ViT [17] proposed by Dosovitskiy et al. is the first milestone of the Transformer-based foundation model. It splits the input image into token representations and introduces a class token before feeding it into the Transformer layers for classification. M. Cornia et al. [39] are pioneers in applying the Transformer model to the image captioning task. S. Herdade et al. [40] integrate spatial information into image captioning by fusing geometric attention weights into the attention mechanism.
After that, the Transformer network is also introduced into the multi-modal community, especially for the vision-language-based pre-training and downstream tasks [41]. Many representative vision-language models are proposed one after another, such as the CLIP [24], ALIGN [42], LXMERT [43], ViLBERT [44]. Inspired by these works, we adopt the Transformer-based pre-trained models to extract the visual and textual representations. More importantly, we formulate the pedestrian attribute recognition task as a sequential text generation problem to better capture the semantic relations between various attributes.
II-C Sequence Generation Models
Many high-level tasks involve generating sequential outputs, such as machine translation, image captioning, and speech recognition. These models usually follow the encoder-decoder framework and achieve this target via recurrent neural networks (RNN) or Transformer decoders. More in detail, the LSTM (Long Short-Term Memory) [33] is widely used for temporal information processing. Chen et al. [45] investigate the impact of different label orders when utilizing LSTMs for multi-label recognition tasks. Attention mechanisms are employed to address the issue of encoder-decoder models when dealing with fixed-length vectors. Lu et al. [46] introduce the concept of a visual sentinel which enables the model to dynamically determine whether to prioritize the visual image or the linguistic rule during different stages of the decoding process. Chen et al. [47] formulate the visual tracking as a sequence generation problem, which predicts object bounding boxes in an auto-regressive fashion. Chen et al. [48] cast object detection as a language modeling task conditioned on the observed pixel inputs and propose a simple and generic framework for object detection termed Pix2Seq. Li et al. [49] propose the BLIP-2 which bootstraps vision-language pre-training using a lightweight Querying Transformer. This work demonstrates that image-to-text generation can follow the introductions of natural language. In this work, we formulate pedestrian attribute recognition as a sequence generation problem instead of the standard multi-label classification problem.
III Methodology
In this section, we will first give an introduction to the existing PAR frameworks and our newly proposed ones. Then, we will introduce the overview of our proposed SequencePAR framework and the input representations of our model, including pedestrian images and attribute phrases. The vision and text encoding process will be provided in detail. After that, we focus on the attribute generation using the Transformer decoder network. Finally, we will introduce the loss function used in our training phase.
Figure 2: An overview of our proposed SequencePAR framework. Given the pedestrian image, we first partition it into non-overlapping patches and extract the visual tokens using pre-trained big models. Unlike previous works that directly regress the attribute scores utilizing a set of dense layers, in this work, we design a new generative paradigm that describes attributes sequentially. It is mainly achieved by a Transformer decoder, which takes the attribute query tokens and pedestrian image features as the input. More in detail, the masked multi-head attention layers are proposed to process the attribute query tokens. Note that, the masking mechanism ensures the model relies on contextual understanding for attribute prediction, preventing opportunistic memorization of subsequent attributes while predicting the current one. We adopt the multi-head cross-attention to fuse the visual and processed query tokens for attribute sequence generation.
III-A Problem Formulation
Pedestrian attribute recognition targets predicting human attributes $\mathcal{A}_{i}$, $i=\{1,2,...,N\}$, from a pre-defined attribute set $\mathcal{A}_{set}=\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{M}\}$, $N\leq M$, based on a given pedestrian image $\mathcal{I}$. Existing deep learning-based attribute recognition models usually treat this task as a multi-label classification or multi-task learning problem [18, 15]. They usually adopt a backbone network to extract the deep features of the human image and learn to recognize the attributes using multi-category classifiers (MCC), i.e.,
$\mathcal{A}_{i}=MCC(VB(\mathcal{I})),$ (1)
where $VB$ denotes the visual backbone. Some researchers have proposed to recognize human attributes by combining pedestrian image features and attribute text representations.
$\mathcal{A}_{i}=MCC(MMFormer(VB(\mathcal{I}),TB(\mathcal{A}_{set}))),$ (2)
where $MMFormer$ denotes the multi-modal Transformer, $TB$ and $\mathcal{A}_{set}$ denotes the text backbone and raw attribute set, respectively. Different from previous works, in this paper, we propose a new sequence generation framework for pedestrian attribute recognition,
$\mathcal{A}_{i}=SGNet(VB(\mathcal{I}),TB(\mathcal{A}_{set})),$ (3)
where $SGNet$ is short for sequence generation network.
III-B Overview
In this work, we formulate the pedestrian attribute recognition task as a sequence generation problem similar to image captioning and machine translation. Given the pedestrian image, we first partition and project it into token representations as the input of the Transformer network. To capture the high-level semantic information of attributes, we also take the raw attributes as the input and utilize the embedded attribute features as the attribute query tokens. The attribute query tokens are processed by masked self-attention layers and further fused with pedestrian image features via cross-attention layers. Such a joint processing of image and text features facilitates the effective integration of multi-modal data for attribute prediction. Finally, the resulting features pass through a linear layer followed by a Softmax function to generate predicted probabilities for each attribute. Therefore, we can get the final attributes via rearrangement. An overview of our proposed SequencePAR framework can be found in Fig. 2. In the following subsections, we will focus on each detailed procedure to help the readers better understand our proposed SequencePAR model.
III-C Network Architecture
Image Encoding Network.  Given a pedestrian attribute recognition dataset $\mathcal{D}=\{(X_{i},A_{i})|i=\{0,1,2,...,N\}\}$, where $X_{i}$ and $A_{i}=\{a_{i}^{1},a_{i}^{2},...,a_{i}^{M}\}$ represents pedestrian images and corresponding attribute labels, respectively. $N$ is the number of pedestrian images in the dataset and $M$ represents the number of labeled attributes. We first split the image into $256$ non-overlapping patches $P_{i}\in\mathbb{R}^{14\times 14}$, then, project them into token representation $T_{i}\in\mathbb{R}^{256\times 1024}$ using one convolutional layer (kernel size 14 $\times$ 14). Following the standard vision Transformer networks, we also introduce the position encoding $PE\in\mathbb{R}^{257\times 1024}$ to capture the spatial information of split patches. The token features and position features are added together as the input of the vision Transformer network. In this paper, we adopt the pre-trained vision encoder of multi-modal big model CLIP [24] (ViT-L/14 is adopted) to extract the feature representations $\hat{T}_{I}\in\mathbb{R}^{257\times 768}$. The core components of ViT-L/14 are the Transformer layers and each layer consists of multi-head self-attention layers, normalization layers, feed-forward layers, and skip connections. Among them, the detailed procedure of multi-head self-attention is the self-attention operation, which can be written as:
$SA(Q,K,V)=Softmax(\frac{QK^{T}}{\sqrt{d}}V),$ (4)
where $T$ denotes the transpose operation, $d$ is the dimension of input tokens.
Attribute Encoding Network.  To help the models better understand the pedestrian attributes that need to be classified, in this work, we take the raw attribute phases $\mathcal{A}_{set}=\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{M}\}$ as the input. To be specific, we expand and split each attribute phase $\mathcal{A}_{i}$ into text representations with the help of prompt engineering. For example, the attribute "age $\leq$ 40" is firstly processed to "age less than 40", then, transformed into a sentence "the age of this pedestrian is less than 40 years old". After that, we obtain the text tokens $\mathcal{E}_{set}=\{\mathcal{E}_{1},\mathcal{E}_{2},...,\mathcal{E}_{M}\}$ using word embedding. Then, we adopt the text encoder of CLIP model [24] to learn the high-level semantic features of human attributes. Therefore, we can get the attribute query tokens $\{\mathcal{Q}_{1},\mathcal{Q}_{2},...,\mathcal{Q}_{M}\}$, which learns the concept of each attribute well. We believe this will make our PAR model better understand these attributes and predict the human attributes in a more accurate way.
Sequence Generation Network.  In this work, we formulate the pedestrian attribute recognition task as a novel sequence generation problem. Given the attribute query tokens $\mathcal{Q}_{i},i\in\{1,2,...,M\}$, we first introduce the position encoding tokens and add these tokens as the input of the sequence generation network. Specifically, the query tokens are treated as Key (K), Value (V), and Query (Q) and feed into the masked multi-head attention module. It receives the predicted attributes from the preceding block and adopts the casual mask to ensure that the predicted pedestrian attributes only depend on its previous sequence cues. As shown in Fig. 2, the output attribute $\mathcal{A}_{i}$ only attends to the attributes less than $i$, with the guidance of the attention mask. Then, the normalization layers and skip connections are adopted to process the obtained features and the output will be treated as the query input of the subsequent multi-head self-attention layer. The visual tokens are used as its key and value inputs. Then, the normalization and feed-forward layers are utilized to further enhance the decoded features $U$.
Post-Processing.  After obtaining the decoded features, we propose to map them into a probability distribution of all attributes defined in each attribute recognition dataset using linear layer and Softmax operator, i.e.,
$[p_{1},p_{2},...,p_{M}]=Softmax(Linear(U)).$ (5)
Then, we transform the index information into its corresponding pedestrian attributes and get the final predictions using the re-arrangement operation. For example, we remove the repeated attributes and special padding tokens.
Loss Function.  The Negative Log Likelihood (NLL) loss function is adopted to optimize our network which can be written as:
$Loss=-\frac{1}{N}\sum_{i=1}^{N}(\sum_{j=1}^{M}w_{j}*y_{ij}*p_{ij})$ (6)
where $N$ and $M$ represent the number of input samples and attributes, respectively. $y_{ij}$ is the label of the corresponding attribute of the sample, $p_{ij}$ is the value of the linear layer output after log_softmax, and $w_{j}=e^{|y_{ij}-r_{j}|}$ is the sampling weight corresponding to the class used for balancing, $r_{j}$ represents the class sampling weight.
IV Experiments
In this section, we will first introduce the datasets and evaluation metric in subsection IV-A. The implementation details are given in subsection IV-B. After that, we report our recognition results and compare them with other state-of-the-art algorithms in subsection IV-C. Then, we conduct extensive studies on our newly proposed SequencePAR framework in subsection IV-D. The visualization is also provided in subsection IV-E to help the readers better understand our model. Then, we discuss the difference between our model and existing recurrent neural network-based PAR models and also the limitation analysis in subsection IV-F and IV-G, respectively.
IV-A Datasets and Evaluation Metric
In this paper, we evaluate our framework on six pedestrian attribute datasets, including PETA, RAPv1, RAPv2, PA100K, PETA-ZS, and RAP-ZS. Note that, the last two datasets are split based on zero-shot setting, i.e., no pedestrian images share the same identity in the training and testing subset. A brief introduction to these datasets is given below.
-- PETA dataset [50] contains 19,000 outdoor or indoor pedestrian images and 61 binary attributes. These images are split into training, validation, and testing subset, which contains 9500, 1900, and 7600 images, respectively. In our experiments, we select 35 pedestrian attributes by following the work [50].
-- RAPv1 dataset [51] contains 41,585 pedestrian images and 69 binary attributes, where 33,268 images are used for training. Usually, in current pedestrian attribute recognition algorithms, 51 attributes are selected for training and evaluation.
-- RAPv2 dataset [52] has 84,928 pedestrian images and 69 binary attributes, where 67,943 were used for training. We select 54 attributes for the training and evaluation of our model.
-- PA100K dataset [29] is the largest pedestrian attribute recognition dataset which contains 100,000 pedestrian images, and 26 binary attributes. In our experiments, we split them into a training and validation set which contains 90,000 images, and a testing subset with the rest 10,000 images.
-- PETA-ZS dataset is proposed by Jia et al. based on PETA [50] dataset by following the zero-shot protocol. The training, validation, and testing subset contains 11241, 3826, and 3933 samples. 35 common attributes are adopted for our experiments by following Jia et al. [53].
-- RAP-ZS dataset is developed based on RAPv2 dataset and contains 17062, 4628, and 4928 pedestrian images for training, validation, and testing, respectively. No shared personal identity between the training and inference data. In our experiments, we select 53 attributes for the evaluation by following Jia et al. [53].
For the evaluation metric, the Accuracy, Precision, Recall, and F1-measure are adopted for the experimental comparison. Specifically, the instance-based evaluation metric Accuracy can be expressed as:
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN},$ (7)
where TP is predicting the correct positive sample, TN is predicting the correct negative sample, FP is a negative sample of prediction errors, and FN is a positive sample of prediction errors. The formulation of Precision, Recall, and F1 measure can be expressed as:
$Precision=\frac{TP}{TP+FP},~{}~{}~{}~{}Recall=\frac{TP}{TP+FN},$ (8)
$F1=\frac{2\times Precision\times Recall}{Precision+Recall}.$ (9)
TABLE I: Comparison with state-of-the-art methods on PETA and PA100K datasets. The first and second are shown in red and blue, respectively. "-" means this indicator is not available.
Methods Ref Backbone PETA PA100K
Accuracy Precision Recall F1 Accuracy Precision Recall F1
DeepMAR [54] ACPR 2015 CaffeNet 75.07 83.68 83.14 83.41 70.39 82.24 80.42 81.32
HPNet [29] ICCV 2017 Inception 76.13 84.92 83.24 84.07 72.19 82.97 82.09 82.53
JRL [15] ICCV 2017 AlexNet - 82.55 82.12 82.02 - - - -
GRL [55] IJCAI 2018 Inception-V3 - 84.34 88.82 86.51 - - - -
MsVAA [56] ECCV 2018 ResNet101 78.56 86.79 86.12 86.46 - - - -
RA [57] AAAI 2019 Inception-V3 - 84.69 88.51 86.56 - - - -
VRKD [58] IJCAI 2019 ResNet50 80.95 88.37 87.47 87.91 78.49 88.42 86.08 87.24
AAP [59] IJCAI 2019 ResNet50 79.95 87.58 87.73 87.65 78.30 89.49 84.36 86.85
VAC [60] CVPR 2019 ResNet50 - - - - 79.44 88.97 86.26 87.59
ALM [61] ICCV 2019 BN-Inception 79.52 85.65 88.09 86.85 77.08 84.24 88.84 86.46
JLAC [14] AAAI 2020 ResNet50 80.38 87.81 87.09 87.50 79.47 87.45 87.77 87.61
SSCsoft [62] ICCV 2021 ResNet50 78.95 86.02 87.12 86.99 78.89 85.98 89.10 86.87
IAA-Caps [63] PR 2022 OSNet 78.04 86.08 85.80 85.64 80.31 88.36 88.01 87.80
MCFL [64] NCA 2022 ResNet50 78.89 84.57 88.84 86.65 77.80 85.11 88.20 86.62
DRFormer [22] NC 2022 ViT-B/16 81.30 85.68 91.08 88.30 80.27 87.60 88.49 88.04
VAC-Combine [65] IJCV 2022 ResNet50 - - - - 80.66 88.72 88.10 88.41
DAFL [66] AAAI 2022 ResNet50 78.88 85.78 87.03 86.40 80.13 87.01 89.19 88.09
CGCN [13] TMM 2022 ResNet50 79.30 83.97 89.38 86.59 - - - -
CAS-SAL-FR [67] IJCV 2022 ResNet50 79.93 87.03 87.33 87.18 79.64 86.81 87.79 85.18
VTB [18] TCSVT 2022 ViT-B/16 79.60 86.76 87.17 86.71 80.89 87.88 89.30 88.21
FEMDAR [68] SPIC 2023 ResNet50 78.45 86.79 85.69 85.90 79.65 87.99 87.45 87.32
EALC [69] NC 2023 EfficientNet-B4 81.71 88.58 88.23 88.40 80.27 87.32 88.98 88.14
APTM [8] ACM MM 2023 Swin-B - - - - 80.17 88.31 87.84 88.07
PARFormer-L [21] TCSVT 2023 Swin-L 82.86 88.06 91.98 89.06 81.13 88.09 91.67 88.52
DFDT [70] EAAI 2023 Swin-B 81.17 87.44 88.96 88.19 81.24 88.02 89.48 88.74
OAGCN [71] TMM 2023 Swin-B 82.95 88.26 89.10 88.68 80.38 84.55 90.42 87.39
VTB* [18] TCSVT 2022 ViT-L/14 79.59 86.66 87.82 86.97 81.76 87.87 90.67 88.86
SequencePAR - ViT-L/14 84.92 90.44 90.73 90.46 83.94 90.38 90.23 90.10
TABLE II: Comparison with state-of-the-art methods on RAPv1 and RAPv2 datasets. The first and second are shown in red and blue, respectively. "-" means this indicator is not available.
Methods Ref Backbone RAPv1 RAPv2
Accuracy Precision Recall F1 Accuracy Precision Recall F1
DeepMAR [54] ACPR 2015 CaffeNet 62.02 74.92 76.21 75.56 - - - -
HPNet [29] ICCV 2017 Inception 65.39 77.33 78.79 78.05 - - - -
JRL [15] ICCV 2017 AlexNet - 75.08 74.96 74.62 - - - -
GRL [55] IJCAI 2018 Inception-V3 - 77.70 80.90 79.29 - - - -
MsVAA [56] ECCV 2018 ResNet101 - - - - 65.57 77.37 79.17 78.26
RA [57] AAAI 2019 Inception-V3 - 79.45 79.23 79.34 - - - -
VRKD [58] IJCAI 2019 ResNet50 69.79 82.13 80.35 81.23 - - - -
AAP [59] IJCAI 2019 ResNet50 68.37 81.04 80.27 80.65 - - - -
VAC [60] CVPR 2019 ResNet50 - - - - 64.51 75.77 79.43 77.10
ALM [61] ICCV 2019 BN-Inception 68.17 74.71 86.48 80.16 64.79 73.93 82.03 77.77
JLAC [14] AAAI 2020 ResNet50 69.15 79.31 82.40 80.82 64.42 75.69 79.18 77.40
SSCsoft [62] ICCV 2021 ResNet50 68.37 75.05 87.49 80.43 - - - -
IAA-Caps [63] PR 2022 OSNet 68.47 79.56 82.06 80.37 - - - -
MCFL [64] NCA 2022 ResNet50 67.28 73.44 87.75 79.96 - - - -
DRFormer [22] NC 2022 ViT-B/16 70.60 80.12 82.77 81.42 - - - -
VAC-Combine [65] IJCV 2022 ResNet50 70.12 81.56 81.51 81.54 - - - -
DAFL [66] AAAI 2022 ResNet50 68.18 77.41 83.39 80.29 66.70 76.39 82.07 79.13
CGCN [13] TMM 2022 ResNet50 54.40 60.03 83.68 70.49 - - - -
CAS-SAL-FR [67] IJCV 2022 ResNet50 68.59 77.56 83.81 80.56 - - - -
VTB [18] TCSVT 2022 ViT-B/16 69.44 78.28 84.39 80.84 67.48 76.41 83.32 79.35
FEMDAR [68] SPIC 2023 ResNet50 66.88 79.11 79.24 78.76 - - - -
EALC [69] NC 2023 EfficientNet-B4 69.65 79.82 83.61 81.67 - - - -
PARFormer-L [21] TCSVT 2023 Swin-L 69.94 79.63 88.19 81.35 - - - -
DFDT [70] EAAI 2023 Swin-B 70.89 80.36 84.32 82.15 69.30 79.38 82.62 80.97
OAGCN [71] TMM 2023 Swin-B 69.32 78.32 87.29 82.56 - - - -
VTB* [18] TCSVT 2022 ViT-L/14 69.78 78.09 85.21 81.10 67.58 76.19 84.00 79.52
SequencePAR - ViT-L/14 71.47 82.40 82.09 82.05 70.14 81.37 81.22 81.10
TABLE III: Comparison with state-of-the-art methods on PETA-ZS and RAP-ZS datasets.
Methods Ref Backbone PETA-ZS RAP-ZS
Accuracy Precision Recall F1 Accuracy Precision Recall F1
MsVAA[56] ECCV 2018 ResNet101 58.67 74.65 69.42 71.94 62.13 75.67 75.81 75.74
VAC[60] CVPR 2019 ResNet50 57.72 72.05 70.64 70.90 63.25 76.23 76.97 76.12
ALM[61] ICCV 2019 BN-Inception 57.78 69.50 73.69 71.53 63.22 72.96 80.73 76.65
JLAC[14] AAAI 2020 ResNet50 58.66 71.70 72.41 72.05 62.58 73.14 79.20 76.05
Jia et al.[53] - ResNet50 58.19 73.09 70.33 71.68 63.61 76.88 76.62 76.75
MCFL[64] NCA 2022 ResNet50 57.04 68.47 74.35 71.29 63.37 71.21 83.86 77.02
VTB [18] TCSVT 2022 ViT-B/16 60.50 73.29 74.40 73.38 64.73 74.93 80.85 77.35
VTB*[18] TCSVT 2022 ViT-L/14 63.12 74.77 77.24 75.50 68.34 76.81 84.51 80.07
SequencePAR - ViT-L/14 66.70 78.75 78.52 78.40 70.28 82.13 80.55 81.14
IV-B Implementation Details
In our experiments, we adopt the ViT-L/14 version of the CLIP [24] model as our feature extractor. The vision encoder of our model is a 24-layer vision Transformer [17] network and the dimension of the hidden layer is 1024. Our sequence generation network is a 6-layer Transformer Decoder, each with eight attention headers. The dimension of the hidden layer is 768. Considering the pedestrian images are all long strip shapes, however, the resolution of input images for the CLIP is $224\times 224$. To adapt to the input resolution, we padded the raw pedestrian images using black pixels into a square shape and resized them into $224\times 224$. Note that, the randomly crop and flip operations are also adopted for data augmentation.
During the training phase, we utilize the ground truth as the input and predict the attributes in a parallel manner, which can reduce the cumulative error significantly and make our model converge faster. For the model inference, we adopt a step-by-step generation approach that uses the output predicted in the previous step and historical output as the inputs of our network. For the detailed parameters, we set the learning rate as 1e-5. We train our model for 50 epochs with the Adam optimizer [72] and set the batch size as 32. Our source code is implemented using Python and the deep learning framework PyTorch [73]. The experiments are conducted based on a server with GPU A100. More details can be found in our source code.
IV-C Comparison with Other SOTA Models
In this section, we will report our recognition results on all six datasets and compare them with existing state-of-the-art pedestrian attribute recognition algorithms. Note that the results of VTB* are obtained by replacing the backbone network of VTB using the visual encoder of CLIP model.
Results on PETA dataset.  As shown in Table I, our proposed SequencePAR model achieves the best performance on most of the evaluation metrics, i.e., 84.92, 90.44, 90.73, 90.46 on the Accuracy, Precision, Recall, and F1-measure, respectively. Compared with other Transformer-based pedestrian attribute recognition models, we can find that our model exceeds both the VTB (ViT-base/16, 79.60/86.76/87.17/86.71) and VTB* (ViT-L/14, 79.59/86.66/87.82/86.97) by +5.32/+3.68/+3.56/+3.67 and +5.33/+3.78/+2.91/+3.41. Our model also achieves better performance than the DRFormer [22] on the Accuracy, Precision, and F1-measure metric. Therefore, we can draw the conclusion that our model achieves state-of-the-art results on the PETA dataset, which fully validates the superiority of our sequence generation framework for the pedestrian attribute recognition task.
Results on PA100K dataset.  As shown in Table I, we can find that our model SequencePAR achieves 83.94/90.38/90.23/90.10 on the Accuracy/Precision/Recall/F1 metric, respectively. These results are better than the multi-label classification model VTB* (81.76/87.87/90.67/88.86) which also adopts the ViT-L/14 model as the backbone network. These experiments demonstrate that our attribute generation-based scheme works better than the discriminative-based learning framework. Also, our model achieves better results than the compared state-of-the-art recognition models.
Results on RAPv1 dataset.  As shown in Table II, we can find that our model achieves 71.47/82.40/82.09/82.02 on the Accuracy/Precision/Recall/F1 metric, respectively. In contrast, the strong baseline VTB* which is the discriminative method achieves 69.78/78.09/85.21/81.10 on these metrics. Obviously, our results are better than theirs, and also the compared attribute recognition models, including CAS-SAL-FR  [67] (IJCV-2022), DRFormer [22] (NC 2022). These experiments fully validated the effectiveness of our proposed model for pedestrian attribute recognition.
Results on RAPv2 dataset.  As shown in Table II, we can find that the multi-label classification based attribute recognition model VTB* (backbone network ViT-L/14) achieves 67.58/76.19/84.00/79.52 on the Accuracy/Precision/Recall/F1 metric, respectively. Meanwhile, our model achieves 70.14, 81.37, 81.22, 81.10. Therefore, we can draw the conclusion that our model is better than the strong discriminative learning based pedestrian attribute recognition algorithm.
Results on PETA-ZS dataset.  In addition to the aforementioned standard evaluation, in this work, we also conduct experiments on the zero-shot setting based datasets. According to the experimental results reported in Table III, we can find that our model achieves 66.70/78.75/78.52/78.40 on the Accuracy/Precision/Recall/F1 metric, respectively. These results are better than the multi-label classification model VTB* (63.12/74.77/77.24/75.50) which also adopts the ViT-L/14 model as the backbone network.
Results on RAP-ZS dataset.  For the experimental results reported on the RAP-ZS dataset as illustrated in Table III, we can find that the baseline method VTB* which achieves 68.34/76.81/84.51/80.07 on the Accuracy/Precision/Recall/F1 metric, respectively. However, we get better results on these metrics, i.e., 70.28, 82.13, 80.55, 81.14. Based on the experiments on the PETA-ZS and RAP-ZS datasets, we can find that our model also works better in this setting. We believe this may be because we use a large visual-language model, which has better generalization ability. Another key reason is that the generative learning based strategy works better for pedestrian attribute recognition.
IV-D Ablation Study
In this work, we conduct extensive experiments to help the readers better understand the interesting factors related to our proposed SequencePAR model.
Greedy Search or Beam Search for Attribute Generation? For the sequence generation tasks, like machine translation and image/video captioning, the greedy search and beam search are all widely used in the inference phase. The greedy search policy will select and remain the element with the maximum response score. In contrast, the beam search policy will always select and retain a fixed number of candidates (also called beam width $\mathcal{B}$) for each time step, and select the trajectory with the maximum summarised response.
In this part, we test different search policies to check whether they still work for pedestrian attribute generation. Specifically, we can find that the greedy search achieves $84.74,90.38,90.42,90.28$ on the Accuracy, Precision, Recall, and F1 measure, as shown in Table IV. When beam search is adopted, the recognition performance on the PETA dataset is similar to the greedy search, however, no improvements are obtained. Therefore, we adopt the simple greedy search scheme for other experiments in this work.
TABLE IV: Compare the greedy search and beam search with different beam widths. The best results are highlighted in bold.
Beam Width PETA
Accuracy Precision Recall F1
1 84.92 90.44 90.73 90.46
3 84.90 90.33 90.79 90.45
5 84.89 90.36 90.76 90.44
10 84.84 90.27 90.80 90.41
Results using Different Layers of Decoders. As shown in Table V, we also conduct experiments on the PETA dataset to check the influence of different layers used in our decoder network. To be specific, we set the layers as 1, 3, 6, 9, and 12. From Table V, we can find that the F1 measures are 89.80, 89.68, 90.46, 89.60, and 89.76, which are relatively stable. When 6 decoding layers are used, we get the best recognition performance on the PETA dataset. Therefore, we default set six layers as our decoder network when conducting other experiments in this paper.
TABLE V: Comparison of Different Decoding Layers for PAR on PETA dataset. The best results are highlighted in bold.
Decoder Layers PETA
Accuracy Precision Recall F1
1 83.96 89.90 89.95 89.80
3 83.84 89.82 89.78 89.68
6 84.92 90.44 90.73 90.46
9 83.72 89.67 89.78 89.60
12 83.90 89.93 89.84 89.76
IV-E Visualization
In this section, we will give a visualization of the similarity matrix of learned pedestrian attributes. Then, we show the predicted pedestrian attributes using our proposed SequencePAR model.
Figure 3: Visualization of the similarity matrix of pedestrian attributes on the PETA dataset obtained by our SequencePAR model.
Figure 4: Average accuracy of attributes on the PETA dataset.
Similarity Matrix.  As shown in Fig 3, we show the cosine similarity between various pedestrian attributes defined in the PETA datasets. We can find that some attributes are relatively independent, such as long hair, upper casual, upper formal, and upper jacket, as shown in the red rectangle. For attributes like age, shoes, attributes defined in the head region are highly correlated, as shown in blue rectangles. In our future works, we will consider learning the human attributes from the highly correlated groups.
Average Accuracy of Pedestrian Attribute.  As shown in Fig. 4, we compare the average accuracy of our method with VTB [18] on 20 pedestrian attributes on the PETA dataset. Our proposed method performs better than VTB [18] on most attributes, and some of the characteristics show significant improvement. Examples of these enhancements include “head sunglasses", “upper jackets", and “shoe sandals", which can improve by up to +12%. These results demonstrate the effectiveness of our generative model for pedestrian attribute recognition.
Figure 5: Visualization of the predicted pedestrian attributes using our proposed SequencePAR.
Pedestrian Attribute Recognition Results.  As shown in Fig 5, we give a visualization of predicted attributes of 15 pedestrian images on the PETA dataset. It is evident from our research that our SequencePAR framework accurately predicts attributes such as age, gender, carried items, clothing, etc. These visualizations fully validate the effectiveness of our generative model for the pedestrian attribute recognition problem.
IV-F Difference with Previous RNN-based PAR
Recurrent Neural Networks (RNN) are widely exploited in pedestrian attribute recognition for feature enhancement [55]. To be specific, Zhao et al. [55] propose to mine both the semantic and spatial correlations in attribute groups and predict the attributes in a group-by-group manner using the LSTM network. Wang et al. [15] propose the JRL model which can jointly learn the attribute context and correlation using the LSTM. Wang et al. [74] propose a recurrent memorized-attention module for multi-label image classification which is interpretable and contextualized. It contains a spatial transformer layer to locate attentional regions from the convolutional feature maps and an LSTM network to sequentially predict semantic labeling scores on the located regions. We can find that these works usually adopt RNN/LSTM for feature enhancement between different human parts, attentional regions mining, or correlations mining between human attributes.
In this paper, we propose a new Transformer-based generative prediction framework for pedestrian attribute recognition. The difference between our work and existing ones can be summarized as follows: 1). Existing works usually adopt the RNN/LSTM to model the attribute relations, while we adopt the advanced self-attention based Transformer network to achieve this goal. Therefore, our model performs better on parallelism, long-range dependencies, attention mechanisms, etc. 2). Many existing RNN/LSTM-based PAR models belong to the discriminative models, meanwhile, our proposed SequencePAR is a new generative framework for the pedestrian attribute recognition task. 3). Existing works predict the pedestrian attributes using the RNN/LSTM network which is trained from scratch, in contrast, our SequencePAR can make full use of the off-the-shelf pre-trained Transformer networks (e.g., vision model CLIP [24] and language model BERT [31]). Therefore, our proposed SequencePAR is significantly different from previous recurrent neural network based attribute recognition models. Also, according to the experimental results reported in sub-section IV-C, our model also exceeds the RNN based models on all the compared benchmark datasets.
IV-G Limitation Analysis
Although good performance can be achieved by our proposed SequencePAR model as validated in previous experiments, however, our model still performs poorly in some challenging scenarios. As illustrated in Fig. 6, some of our predicted pedestrian attributes are incorrect which are highlighted in red. It is easy to find that these images usually contain multiple persons or interfering objects which are really hard for the model to discriminate. Therefore, it will be a possible solution to first locate the true pedestrian we need to analyze, and then, predict the pedestrian attributes. We leave this as our future works.
Figure 6: Visualization of some incorrectly predicted pedestrian attributes (highlighted in red) using our proposed SequencePAR.
V Conclusion
In this work, we propose a novel sequence generation paradigm for pedestrian attribute recognition, termed SequencePAR. It is proposed based on the fact that existing pedestrian attribute recognition (PAR) algorithms are following multi-label or multi-task learning. That is to say, their performance heavily depends on the specific classification heads and is easily influenced by imbalanced data or noisy samples. Therefore, we propose the SequencePAR which belongs to the generative models. To be specific, it extracts the pedestrian features using a pre-trained CLIP model and embeds the attribute set into query tokens under the guidance of text prompts. Then, a Transformer decoder is proposed to generate the human attributes by incorporating the visual features and attribute query tokens. The masked multi-head attention layer is introduced into the decoder module to prevent the model from remembering the next attribute while making attribute predictions during training. Extensive experiments on multiple widely used PAR datasets fully validated the effectiveness of our proposed SequencePAR. In our future work, we will consider introducing new main target pedestrian localization modules to eliminate the interference of background pedestrians, as discussed in Section IV-G.